Hey
Im really struggling with my math problems
a motor cyclist makes a trip of 500km. If he had increased his speed by 10 km/h, he could have covered 600km in the same time. What was his original speed?
I need the steps
Thank you

Answer by mananth(16946)

distance = 500
speed be x kmph
...
time = 500/x
...
increase speed by 10 kmph
speed = x+10
distance = 600
time = 600/x+10
...
both times are equal
600/x+10=500/x
600x=500(x+10)
600x=500x+5000
600x-500x=5000
100x=5000
x= 50 kmph
..
m.ananth@hotmail.ca